# Admissibility Breakdown in High-Dimensional Sparse Regression with $\ell_{1}$ Regularization

### v1

@@ Abstract
The choice of the tuning parameter in the Lasso is central to its statistical performance in high-dimensional linear regression. Classical consistency theory identifies the rate of the Lasso tuning parameter, and numerous studies have established non-asymptotic guarantees. Nevertheless, the question of optimal tuning within a non-asymptotic framework has not yet been fully resolved. We establish tuning criteria above which the Lasso becomes inadmissible under mean squared prediction error. More specifically, we establish thresholds showing that certain classical tuning choices yield Lasso estimators strictly dominated by a simple Lasso–Ridge refinement. We also address how the structure of the design matrix and the noise vector influences the inadmissibility phenomenon.
@@ 1 Introduction
The Lasso (Tibshirani, 1996) is one of the most widely used methods for high-dimensional linear regression. By combining estimation and variable selection through an $\ell_{1}$ penalty, it leads to computationally efficient procedures with solid theoretical properties. A large body of work has established oracle inequalities, prediction error bounds, and sparsity recovery guarantees under various assumptions on the design matrix; see, for example, (van de Geer, 2016) and the references therein. In this literature, it is standard to choose the tuning parameter at the rate $\lambda\asymp\sigma\sqrt{\log(p)/n}$, which ensures favorable prediction behavior in high-dimensional settings.
More recently, several studies have suggested that the choice of the tuning parameter can be more delicate, especially when the design exhibits strong correlations. For instance, (van de Geer and Lederer, 2013; Hebiri and Lederer, 2013; Dalalyan et al., 2017) show that tuning parameters substantially smaller than the classical choice may yield sharp oracle inequalities and, in some cases, improved prediction performance. Related results in (Hebiri and Lederer, 2013) indicate that stronger correlations in the design matrix call for smaller tuning parameters to obtain sharper bounds. In a different direction, (Bellec, 2018) introduced the notion of a “noise barrier,” showing that the minimal penalty level may depend on intrinsic quantities such as sparsity, leading to refined scales of order $\lambda\asymp\sigma\sqrt{\log(p/k)/n}$ rather than $\sigma\sqrt{\log(p)/n}$. Taken together, these works provide theoretical support for the empirical observation that tuning parameters smaller than the classical rate can be advantageous in certain regimes. However, these results are primarily expressed in terms of refined risk bounds and do not directly address whether the Lasso equipped with the traditional tuning parameter can be improved upon in a uniform sense.
In this paper, we provide a decision-theoretic perspective on Lasso tuning and prediction improvement. Rather than focusing on refined prediction bounds with new tuning choices, we ask whether the Lasso, for a given tuning parameter, can be dominated in mean squared prediction error by a suitably constructed refinement, as proposed in a previous arXiv preprint (Liu, 2025). We show that, over a broad range of tuning values, including the standard Lasso tuning, the Lasso estimator is dominated in prediction risk by an Lasso–Ridge refinement. Although the Lasso serves as our primary example, the analysis extends, to some extent, to other sparse estimators induced by $\ell_{1}$ regularization. Our results provide a complementary explanation for the empirical finding that tuning parameters smaller than the standard theoretical choice often yield superior prediction performance. This perspective also highlights a distinction between rate-based guarantees and uniform risk improvement, and clarifies how tuning choices interact with sparsity and design structure.
The main contributions of this paper can be summarized as follows:
@@
    We introduce a framework for comparing prediction risks that allows direct assessment of whether a given Lasso tuning choice can be improved upon.
@@
    We establish criteria for Lasso tuning parameters beyond which the Lasso is dominated in mean squared prediction error, showing that a range containing commonly used tuning parameters can lead to suboptimal performance.
@@
    We provide insight into how intrinsic structural features, such as sparsity and properties of the design matrix, affect the inadmissibility phenomenon.
@@ 2.1 Notation and Model Assumptions
For any $S\subseteq[p]$, $A_{S}$ denotes the submatrix consisting of the columns of $A$ indexed by $S$, and $\beta_{S}$ denotes the corresponding subvector of $\beta$. We adopt the standard conventions for zero-dimensional vectors and matrices. In particular, a $0\times n$ matrix and a $0\times 1$ vector are well defined, and the $0\times 0$ matrix is treated as a valid square matrix whose inverse is itself. Under these conventions, matrix multiplication involving zero-dimensional objects is well defined and dimensionally consistent. We work with the linear model
$y=X\beta^{0}+\varepsilon,$
where $y\in\mathbb{R}^{n}$, $X\in\mathbb{R}^{n\times p}$, and $\beta^{0}\in\mathbb{R}^{p}$. The noise vector $\varepsilon\in\mathbb{R}^{n}$ is arbitrary and no specific distributional assumption is imposed unless stated otherwise. Columns of $X$ are normalized so that $\|X_{j}\|_{2}^{2}=n$ for all $j=1,\dots,p$.
@@ 2.2 Lasso and Ridge: Definitions and Basic Properties
Definition 2.1.
The Lasso estimator is noted as $\hat{\beta}_{\mathrm{L}}$, with its definition as follows:
$\hat{\beta}_{\mathrm{L}}=\underset{\beta\in\mathbb{R}^{p}}{\arg\min}\left\{\frac{1}{2n}\|y-X\beta\|_{2}^{2}+\lambda_{\mathrm{L}}\|\beta\|_{1}\right\},$
Proposition 2.2.
The KKT conditions for the Lasso estimator follow from subdifferential calculus (van de Geer, 2016). They can be expressed as
$\frac{1}{n}X^{\top}(y-X\hat{\beta}_{\mathrm{L}})=\lambda_{\mathrm{L}}\bar{z},$
where $\bar{z}\in\partial\|\hat{\beta}_{\mathrm{L}}\|_{1}$, with the $i$-th component given by
Also, notice that:
$\frac{1}{n}X_{E}^{\top}(y-X\hat{\beta}_{\mathrm{L}})=\lambda_{\mathrm{L}}s.$
Here, $\mathrm{sign}(\cdot)$ denotes the sign function. In addition, the KKT condition on the active block derives
Assumption 2.3.
The columns of the design matrix $X$ are in general position.
Proposition 2.4 (Uniqueness of the Lasso).
Under Assumption 2.3, the Lasso solution $\hat{\beta}_{\mathrm{L}}$ is unique, with support exactly $E$; that is, $\hat{\beta}_{j}\neq 0$ iff $j\in E$. Also, $\mathrm{null}(X_{E})=\{0\}$. See Tibshirani (2013) for further details.
Remark 2.5.
Assumption 2.3 ensures that the definition of the refined estimator is unambiguous. Throughout this paper, we assume that Assumption 2.3 holds without further mention.
@@
    $s\coloneqq\operatorname{sign}({X_{E}}^{T}(y-X\hat{\beta}_{\mathrm{L}})).$
@@
By the polyhedral lemma (Lee et al., 2016), conditioning on a fixed active set $E=E_{i}$ and sign vector $s=s_{i}$ is equivalent to restricting the random noise vector $\varepsilon$ to lie in a corresponding polyhedral region. Specifically, there exist a matrix $A_{i}$ and a vector $b_{i}$ such that
$\mathbf{1}_{\{E=E_{i},\,s=s_{i}\}}=\mathbf{1}_{\{A_{i}\varepsilon\leq b_{i}\}}=\mathbf{1}_{\{\varepsilon\in\mathcal{P}_{i}\}},$
where $\mathcal{P}_{i}=\{\varepsilon:A_{i}\varepsilon\leq b_{i}\}$ denotes the polyhedron associated with the selection event $\{E=E_{i},s=s_{i}\}$.
Proposition 2.8.
The ridge regression estimator (Hoerl and Kennard, 1970) is defined as
$\hat{\beta}^{\mathrm{r}}=\arg\min_{\beta\in\mathbb{R}^{p}}\left\{\frac{1}{2n}\|y-X\beta\|_{2}^{2}+\frac{\lambda^{\mathrm{r}}}{2}\|\beta\|_{2}^{2}\right\},$
where $\lambda^{\mathrm{r}}>0$ is a regularization parameter. The corresponding Karush–Kuhn–Tucker (KKT) condition is
$-\frac{1}{n}X^{\top}\bigl(y-X\hat{\beta}^{\mathrm{r}}\bigr)+\lambda^{\mathrm{r}}\hat{\beta}^{\mathrm{r}}=0,$
which yields the closed-form expression
$\hat{\beta}^{\mathrm{r}}=(X^{\top}X+n\lambda^{\mathrm{r}}I_{p})^{-1}(X^{\top}y).$
@@ 2.3 The Lasso–Ridge Estimator and Related Definitions
A Lasso-Ridge estimator could be defined as:
@@
Let $\hat{\beta}_{\mathrm{L}}$ and $\hat{\beta}_{\mathrm{R}}$ Lasso and Lasso–Ridge estimators of $\beta_{0}$. We define the prediction error gap:
@@ 3 Main Results
$\mathrm{DMSE}$ $\geq\sup_{T_{0}\subseteq\{1,\dots,p\}}\Bigl\{\frac{2\lambda_{\mathrm{L}}}{c}(\frac{c(2c+1)}{2(c+1)^{2}}\lambda_{\mathrm{L}}\mathbb{P}(E\neq\emptyset)-\frac{1}{n}\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}]$
$-\frac{1}{n}\sqrt{\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}^{2}]}\sqrt{\mathbb{P}(E\not\subseteq T_{0})})\Bigr\}.$
@@
where $P_{X_{E}}$ denotes the orthogonal projection onto $\operatorname{col}(X_{E})$ and $P_{X_{E}}^{\perp}=I_{|E|}-P_{X_{E}}$ is the projection onto its orthogonal complement. Inserting this rewritten form into equality (1) and rearranging, we obtain
@@
$\langle P_{X_{E}}^{\perp}X\beta_{0},X_{E}\hat{\delta}_{E}\rangle$ $=\hat{\delta}_{E}^{\top}X_{E}^{\top}(I_{|E|}-X_{E}(X_{E}^{\top}X_{E})^{-1}X_{E}^{\top})X\beta_{0}$
@@
$=\hat{\delta}_{E}\mathbf{0}_{|E|}X\beta_{0}$
@@
$\Sigma_{n,E}=P\operatorname{diag}(\mu_{1}+,\dots,\mu_{|E|})P^{\top}$
where $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{|E|}$ denote the ordered eigenvalues of $\Sigma_{n,E}=X_{E}^{\top}X_{E}/n$ and $P$ is the is the corresponding eigenvector matrix. According to Lemma A.3 in the Appendix, we obtain the following closed-form representation of $\hat{\delta}_{E}$:
@@
$=\lambda_{\mathrm{L}}^{2}s^{\top}(\lambda_{\mathrm{R}}\Sigma_{n,E}+\lambda_{\mathrm{R}}I_{|E|})^{-1}\Sigma_{n,E}(\lambda_{\mathrm{R}}\Sigma_{n,E}+\lambda_{\mathrm{R}}I_{|E|})^{-1}s$
@@
Recall that we choose the tuning parameter $\lambda_{\mathrm{R}}=c|E|$ where $c>2$. Under this choice, we derive the following inequalities:
@@
Next, we consider the remaining part of equation (3). From Lemma A.3 in the appendix, we know the following equation holds when $\lambda_{\mathrm{R}}>2|E|$:
@@
It follows from the dual-norm inequality that
@@
Next, we analysis the term $\mathbb{E}[\|X_{E}^{\top}\varepsilon\|_{\infty}]$ in the following way:
@@
$\geq\sup_{T_{0}\subseteq\{1,\dots,p\}}\Bigl\{\frac{2\lambda_{\mathrm{L}}}{c}(\frac{c(2c+1)}{2(c+1)^{2}}\lambda_{\mathrm{L}}\mathbb{P}(E\neq\emptyset)-\frac{1}{n}\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}]$
$-\frac{1}{n}\sqrt{\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}^{2}]}\sqrt{\mathbb{P}(E\not\subseteq T_{0})})\Bigr\}.$
Lemma 3.2.
Suppose that $\varepsilon\sim N(0,\sigma^{2}I_{n})$ and $\|X_{j}\|_{2}^{2}=n$ for all $j$. Then $\mathbb{E}[\|X^{\top}\varepsilon/n\|_{\infty}^{2}]\leq\sigma^{2}(2\log(4p)+1)/n.$ See Lemma A.1 for a proof.
@@
Let $S_{0}=\{j:\beta_{0,j}\neq 0\}$ denote the true support. We assume the existence of a deterministic index set $T_{0}$ such that
$\mathbb{P}(E\subseteq T_{0})\geq 1-\delta$
for a fixed error probability $0<\delta<1$. When $T_{0}=S_{0}$, $E\subseteq S_{0}$ corresponds to the event that the Lasso has no false alarms, which occurs with high probability under mild conditions (van de Geer, 2016, Section 4.7, p. 51). Next, considering the case where $T_{0}=\{1,\dots,p\}$, we obtain the following corollary:
@@
Consider our Lasso-Ridge method. Select the tuning parameter of the second step as $\lambda_{\mathrm{R}}=c|E|$, where $c>2$, then we have:
@@
Corollary 3.5.
Suppose the penalty level $\lambda_{\mathrm{L}}$ of the Lasso estimator satisfies
@@
In particular, the Lasso cannot achieve the optimal prediction risk whenever $\lambda_{\mathrm{L}}$ slightly exceeds the the expectation of the random maximum.
Remark 3.6.
For further discussions, we first recall the standard bound for $\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}]$. We then discuss how the structure of the design matrix $X$ critically influences the behavior of this Gaussian maximum.
@@
For more details, see (van de Geer, 2016, Section 17.4, p. 243).
Lemma 3.8 (Generic chaining bound).
Let $(T,d)$ be a metric space. If a centered process $\{X_{t}:t\in T\}$ has sub-Gaussian increments with respect to $d$, then there exists an absolute constant $C>0$ such that
$\mathbb{E}\sup_{t\in T}X_{t}\;\leq\;C\,\gamma_{2}(T,d),$
where $\gamma_{2}(T,d)$ is the generic chaining functional
$\gamma_{2}(T,d)=\inf_{\{T_{k}\}}\sup_{t\in T}\sum_{k\geq 0}2^{k/2}\,d(t,T_{k}),$
the infimum being over admissible sequences $\{T_{k}\}$, that is, $|T_{0}|=1$ and $|T_{k}|\leq 2^{2^{k}}$. For more details, see (Vershynin, 2018, Theorem 8.5.3, p. 208).
Lemma 3.9 (Two-step bound under column normalization and clustering).
Let $X=[X_{1},\dots,X_{p}]\in\mathbb{R}^{n\times p}$ satisfy $\|X_{j}\|_{2}=\sqrt{n}$ for all $j$. Assume that the index set $\{1,\dots,p\}$ is partitioned into $K$ clusters $\{\mathcal{C}_{\ell}\}_{\ell=1}^{K}$, and for each cluster fix a representative $r_{\ell}\in\mathcal{C}_{\ell}$. Write $r(j)$ for the representative of the cluster containing $j$, and suppose the clusters are $\delta$-tight in the sense that
$\|X_{j}-X_{r(j)}\|_{2}\leq\delta\sqrt{n},\qquad\forall j,$

### v2

# On the Prediction Suboptimality of the Lasso

@@ Abstract
The choice of the tuning parameter in the Lasso is central to its statistical performance in high-dimensional linear regression. Classical consistency theory characterizes the rate of the Lasso tuning parameter, and numerous works provide non-asymptotic guarantees. However, the optimal choice of tuning within a fully non-asymptotic framework remains incompletely understood. In this work, we study tuning regimes under which the Lasso exhibits suboptimal prediction performance, in the sense that it is dominated in mean squared prediction error by a simple refinement. We further examine how structural factors in the design matrix influence the suboptimality phenomenon, and discuss extensions to other estimators and more general noise structures.
@@ 1 Introduction
Linear regression is a simple yet powerful framework for high-dimensional modeling, valued for its interpretability and its transparent representation of signal and noise. The Lasso (Tibshirani, 1996) is one of the most widely used methods for high-dimensional linear regression. By combining estimation and variable selection through an $\ell_{1}$ penalty, it yields computationally efficient estimators with well-developed theoretical guarantees. A large body of work has established oracle inequalities, prediction error bounds, and sparsity recovery results under various assumptions on the design matrix; see, for example, (van de Geer, 2016) and the references therein. Within this literature, it is common to select the tuning parameter at the universal scale $\sigma\sqrt{(\log p)/n}$, which ensures favorable asymptotic and non-asymptotic guarantees in high-dimensional settings.
More recently, it has been observed that the choice of the tuning parameter can be more delicate, particularly when the design matrix exhibits strong correlations. For instance, related results in (van de Geer and Lederer, 2013; Hebiri and Lederer, 2013) indicate that stronger correlations in the design call for smaller tuning parameters in order to obtain tighter guarantees. In a related direction, (Dalalyan et al., 2017) consider new measures of design correlation and establish refined Lasso oracle inequalities within this framework. From a different perspective, (Bellec et al., 2018) show that adaptively chosen tuning parameters for the Lasso can yield sharp oracle inequalities and improved minimax prediction rates under suitable design conditions. Taken together, these results suggest that tuning parameters below the universal Lasso scale of order $\sigma\sqrt{(\log p)/n}$ may be advantageous in certain regimes. These works primarily focus on refined risk upper bounds and therefore do not fully address this aspect of the problem, and we are interested in whether the presence of $\ell_{1}$-regularization inherently induces a broad range of tuning parameters that lead to suboptimal prediction.
In this paper, we provide new insights into the prediction and tuning of $\ell_{1}$-regularized high-dimensional regression from the perspective of prediction risk dominance within a non-asymptotic framework. Rather than proposing new tuning rules for refined upper bounds, we characterize the range of tuning parameters for which prediction suboptimality occurs and establish lower bounds on the prediction mean squared error difference. Specifically, for the Lasso, when the tuning parameter exceeds approximately $\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}/n],$ where $T_{0}$ denotes a high-probability containment set of the selected support of the Lasso, it can be dominated in prediction mean squared error by a suitably constructed refinement, as suggested in (Liu, 2025). Notably, the lower bound is explicit, with no hidden constants, and the resulting dominance range can include tuning levels at the standard theoretical scale $\lambda\asymp\sigma\sqrt{(\log p)/n}$.
Although the Lasso serves as the primary example, the analysis may extend to a broader class of $\ell_{1}$-regularized estimators and to more general noise structures. The results provide a theoretical explanation for the empirical observation that smaller tuning parameters can yield improved prediction performance, while clarifying the distinction between rate-optimality and finite-sample risk dominance. The main contributions are as follows:
@@
    A framework for direct comparison of prediction risks, allowing one to assess whether a given Lasso tuning choice admits improvement via a simple refinement.
@@
    Explicit range conditions on the tuning parameter under which the Lasso is dominated in mean squared prediction error.
@@
    A characterization of how structural features, including sparsity and design correlations, shape the regimes of prediction risk dominance, together with several extensions.
@@ 2.1 Notation and Model Assumptions
For any $S\subseteq[p]$, $A_{S}$ denotes the submatrix consisting of the columns of $A$ indexed by $S$, $\beta_{S}$ denotes the corresponding subvector of $\beta$, and $\beta_{-S}$ denotes the subvector of $\beta$ indexed by the complement $S^{c}$. We adopt the standard conventions for zero-dimensional vectors and matrices. In particular, a $0\times n$ matrix and a $0\times 1$ vector are well defined, and the $0\times 0$ matrix is treated as a valid square matrix whose inverse is itself. Under these conventions, matrix multiplication involving zero-dimensional objects is well defined and dimensionally consistent. We work with the linear model
$y=X\beta_{0}+\varepsilon,$
where $y\in\mathbb{R}^{n}$, $X\in\mathbb{R}^{n\times p}$, and $\beta^{0}\in\mathbb{R}^{p}$. The noise vector $\varepsilon\in\mathbb{R}^{n}$ is arbitrary and no specific distributional assumption is imposed unless stated otherwise.
@@ 2.2 Lasso and Ridge: Definitions and Basic Properties
Definition 2.1 (Lasso estimator).
Let $\lambda_{\mathrm{L}}>0$ be a regularization parameter. The Lasso estimator $\hat{\beta}_{\mathrm{L}}\in\mathbb{R}^{p}$ is defined as
$\hat{\beta}_{\mathrm{L}}\in\arg\min_{\beta\in\mathbb{R}^{p}}\left\{\frac{1}{2n}\|y-X\beta\|_{2}^{2}+\lambda_{\mathrm{L}}\|\beta\|_{1}\right\}.$
@@
    $s\coloneqq\operatorname{sign}({X_{E}}^{\top}(y-X\hat{\beta}_{\mathrm{L}})).$
@@
The Karush–Kuhn–Tucker (KKT) conditions for the Lasso estimator can be written as follows; see, for example, (van de Geer, 2016):
$\frac{1}{n}X^{\top}(y-X\hat{\beta}_{\mathrm{L}})=\lambda_{\mathrm{L}}\bar{z},$
where $\bar{z}\in\partial\|\hat{\beta}_{\mathrm{L}}\|_{1}$, whose $i$-th component is given by
Moreover, it follows that
$\frac{1}{n}X_{E}^{\top}(y-X\hat{\beta}_{\mathrm{L}})=\lambda_{\mathrm{L}}s.$
Here, $\mathrm{sign}(\cdot)$ denotes the sign function. In addition, the KKT conditions on the active block imply
$\hat{\beta}_{\mathrm{L},E}=(X_{E}^{\top}X_{E})^{-1}\!\left(X_{E}^{\top}y-n\lambda_{\mathrm{L}}s\right).$
Here, $(\hat{\beta}_{\mathrm{L}})_{i}$ denotes the $i$-th component of $\hat{\beta}_{\mathrm{L}}$, and for an index set $E$, $\hat{\beta}_{\mathrm{L},E}$ denotes the subvector restricted to $E$. The explicit expression for $\hat{\beta}_{\mathrm{L},E}$ holds when the Lasso solution is unique (see Proposition 2.6).
Assumption 2.4 (Column normalization).
The columns of the design matrix $X\in\mathbb{R}^{n\times p}$ are normalized as
$\|X_{j}\|_{2}^{2}=n,\qquad j=1,\dots,p.$
Assumption 2.5.
The columns of the design matrix $X$ are in general position. More precisely, for any integer $k<\min\{n,p\}$, no $k$-dimensional affine subspace of $\mathbb{R}^{n}$ contains more than $k+1$ elements of the signed set $\{\pm X_{1},\ldots,\pm X_{p}\}$, excluding antipodal pairs. Although the definition may appear somewhat technical, this condition is naturally satisfied when the entries of $X$ are drawn from a continuous distribution.
Proposition 2.6 (Uniqueness of the Lasso).
Under Assumption 2.5, the Lasso solution $\hat{\beta}_{\mathrm{L}}$ is unique, and its support is exactly $E$; that is, $\hat{\beta}_{j}\neq 0$ if and only if $j\in E$. Moreover, $\mathrm{null}(X_{E})=\{0\}$; see (Tibshirani, 2013).
Remark 2.7.
Proposition 2.6 ensures that the definition of the refined estimator is unambiguous. Throughout this paper, we impose Assumptions 2.4 and 2.5 without further mention.
Proposition 2.8 (Ridge regression estimator).
Let $\lambda^{\mathrm{r}}>0$ be a regularization parameter. The ridge regression estimator $\hat{\beta}^{\mathrm{r}}\in\mathbb{R}^{p}$ (Hoerl and Kennard, 1970) is defined by
$\hat{\beta}^{\mathrm{r}}=\arg\min_{\beta\in\mathbb{R}^{p}}\left\{\frac{1}{2n}\|y-X\beta\|_{2}^{2}+\frac{\lambda^{\mathrm{r}}}{2}\|\beta\|_{2}^{2}\right\}.$
The Karush–Kuhn–Tucker (KKT) condition is given by
$\frac{1}{n}X^{\top}\bigl(y-X\hat{\beta}^{\mathrm{r}}\bigr)+\lambda^{\mathrm{r}}\hat{\beta}^{\mathrm{r}}=0,$
which implies the closed-form solution
$\hat{\beta}^{\mathrm{r}}=\bigl(X^{\top}X+n\lambda^{\mathrm{r}}I_{p}\bigr)^{-1}X^{\top}y.$
@@ 2.3 The Lasso–Ridge Estimator and Related Definitions
A Lasso–Ridge estimator with $\lambda_{\mathrm{R}}\geq 0$ is defined as
@@
Let $\hat{\beta}_{\mathrm{L}}$ and $\hat{\beta}_{\mathrm{R}}$ denote the Lasso and Lasso–Ridge estimators of $\beta_{0}$, respectively. We define the prediction error gap:
@@ 3 Main Results
Consider our Lasso–Ridge method. Select the tuning parameter of the second step as $\lambda_{\mathrm{R}}=c|E|$, where $c>2$, then we have
$\mathrm{DMSE}$ $\geq\sup_{T_{0}\subseteq\{1,\dots,p\}}\Bigl\{\frac{2\lambda_{\mathrm{L}}}{c}\Bigl(\frac{c(2c+1)}{2(c+1)^{2}}\lambda_{\mathrm{L}}\mathbb{P}(E\neq\emptyset)-\frac{1}{n}\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}]$
$\hskip 20.00003pt-\frac{1}{n}\sqrt{\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}^{2}]}\sqrt{\mathbb{P}(E\not\subseteq T_{0})}\Bigr)\Bigr\}.$
Here $|E|$ denotes the Lasso selected model size, which is random, so the tuning parameter $\lambda_{\mathrm{R}}$ is data-dependent.
@@
where $P_{X_{E}}$ denotes the orthogonal projection onto $\operatorname{col}(X_{E})$ and $P_{X_{E}}^{\perp}=I_{n}-P_{X_{E}}$ is the projection onto its orthogonal complement. Inserting this rewritten form into equality (1) and rearranging, we obtain
@@
$\langle P_{X_{E}}^{\perp}X\beta_{0},X_{E}\hat{\delta}_{E}\rangle$ $=\hat{\delta}_{E}^{\top}X_{E}^{\top}(I_{n}-X_{E}(X_{E}^{\top}X_{E})^{-1}X_{E}^{\top})X\beta_{0}$
@@
$=\hat{\delta}_{E}\mathbf{0}_{|E|\times n}X\beta_{0}$
@@
$\Sigma_{n,E}=P\operatorname{diag}(\mu_{1},\dots,\mu_{|E|})P^{\top},$
where $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{|E|}\geq 0$ denote the ordered eigenvalues of $\Sigma_{n,E}=X_{E}^{\top}X_{E}/n$ and $P$ is the corresponding eigenvector matrix. According to Lemma A.2 in the Appendix, we obtain the following closed-form representation of $\hat{\delta}_{E}$:
@@
Recall that we choose the tuning parameter $\lambda_{\mathrm{R}}=c|E|$ where $c>2$. Since $s$ is a binary vector and $P$ is orthogonal (hence norm-preserving), the following inequalities hold:
@@
Next, we consider the last remaining term in equation (3). From Lemma A.3 in the appendix, we know the following equation holds when $\lambda_{\mathrm{R}}>2|E|$:
@@
Returning to the remaining term, the dual-norm inequality yields
@@
Next, we analyze the term $\mathbb{E}[\|X_{E}^{\top}\varepsilon\|_{\infty}]$ in the following way:
@@
$\geq\sup_{T_{0}\subseteq\{1,\dots,p\}}\Bigl\{\frac{2\lambda_{\mathrm{L}}}{c}\Bigl(\frac{c(2c+1)}{2(c+1)^{2}}\lambda_{\mathrm{L}}\mathbb{P}(E\neq\emptyset)-\frac{1}{n}\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}]$
$\hskip 20.00003pt-\frac{1}{n}\sqrt{\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}^{2}]}\sqrt{\mathbb{P}(E\not\subseteq T_{0})}\Bigr)\Bigr\}.$
Remark 3.2.
Suppose that $\varepsilon\sim N(0,\sigma^{2}I_{n})$ under the column normalization of $X$. Then $\mathbb{E}[\|X^{\top}\varepsilon/n\|_{\infty}^{2}]\leq\sigma^{2}(2\log(4p)+1)/n.$ See Lemma A.4 in the appendix for a proof. Moreover, $\mathbb{P}(E\neq\emptyset)$, that is, the probability that the Lasso estimator is nonzero, is typically close to one under standard signal and tuning regimes. See Lemma A.5 and Remark A.6 in the appendix for a more detailed discussion.
Remark 3.3 (On the data-dependent choice of $\lambda_{\mathrm{R}}$).
This tuning choice $\lambda_{\mathrm{R}}=c|E|$ is adopted primarily to derive a clean and easily interpretable lower bound for the prediction MSE difference. The constant $c>2$ is mainly technical and can be relaxed under stronger design conditions, for example when the design matrix has bounded column inner products (i.e., weaker mutual coherence or more nearly orthogonal columns). Other choices of $\lambda_{\mathrm{R}}$ are also possible and would lead to alternative formulations, typically involving probability inequalities or less transparent bounds.
@@
Note that $\sqrt{\mathbb{E}[\|X^{\top}\varepsilon/n\|_{\infty}^{2}]}$ has the same Gaussian-max scale as $\mathbb{E}[\|X^{\top}\varepsilon/n\|_{\infty}]$, while the third term in Theorem 3.1 is further attenuated by the factor $\mathbb{P}(E\not\subseteq T_{0})$. This motivates assuming the existence of a deterministic index set $T_{0}$ such that
$\mathbb{P}(E\subseteq T_{0})\geq 1-\eta$
for some fixed $0<\eta<1$, which controls the negative contribution in the lower bound. Let $S_{0}=\{j:\beta_{0,j}\neq 0\}$ denote the true support. When $T_{0}=S_{0}$, the event $E\subseteq S_{0}$ corresponds to the Lasso having no false inclusions, which holds with high probability under mild conditions; see, for example, (Hastie et al., 2015, Theorem 11.3, p. 302).
The result cited above also illustrates that such containment statements are generally weaker than exact variable selection consistency. In particular, support containment may hold without requiring strong signal conditions on $\beta_{0}$, and the irrepresentable conditions can be relaxed when it is considered with respect to a deterministic superset rather than the exact true support. Under such formulations, the probability of containment failure may still decay exponentially with the sample size.
It is worth noting that, instead of relying solely on the Cauchy–Schwarz inequality, one may invoke Hölder’s inequality, which provides additional flexibility in controlling and potentially attenuating the negative contribution. Next, considering the case $T_{0}=\{1,\dots,p\}$, we obtain the following corollaries:
@@
Remark 3.6.
Consider the function
$f_{2}(x)=\frac{x(2x+1)}{(x+1)^{2}},\quad x\geq 0.$
The function $f_{2}(x)$ is monotone increasing, and it converges to $2$ as $x\to\infty$. Suppose the penalty level $\lambda_{\mathrm{L}}$ of the Lasso estimator satisfies
@@
In particular, since $\mathbb{P}(E\neq\emptyset)$ is typically close to one under the considered regimes, the Lasso can be prediction-suboptimal within this refinement class once $\lambda_{\mathrm{L}}$ moderately exceeds the noise-max scale $\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}]/n$.
Remark 3.7 (Noise robustness).
The main results do not rely on independence or Gaussianity of the noise. In particular, martingale difference errors can be accommodated under mild conditions. See the extension part for details. For further discussions, we first recall the standard bound for $\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}]$. We then discuss how the structure of the design matrix $X$ critically influences the behavior of this Gaussian maximum.
@@
For more details, see, for example, (van de Geer, 2016, Section 17.4, p. 243).
Remark 3.9 (Gaussian process interpretation).
Define $Z_{j}:=n^{-1}X_{j}^{\top}\varepsilon$, $j=1,\dots,p$. Then $\{Z_{j}\}_{j=1}^{p}$ is a centered Gaussian process indexed by $I=\{1,\dots,p\}$, with canonical metric
$d(i,j)=(\mathbb{E}|Z_{i}-Z_{j}|^{2})^{1/2}=\frac{\sigma}{n}\|X_{i}-X_{j}\|_{2},\qquad i,j\in I.$
Moreover,
$n^{-1}\|X^{\top}\varepsilon\|_{\infty}=\sup_{j\in I}|Z_{j}|.$
Thus, the stochastic term may be analyzed using Gaussian process tools, with the sub-Gaussian increment structure determined by the metric geometry of the column set. In particular, this representation makes general tools such as the chaining bound stated later naturally applicable.
Lemma 3.10 (Two-step bound under clustering).
Let $X=[X_{1},\dots,X_{p}]\in\mathbb{R}^{n\times p}$. Assume that the index set $\{1,\dots,p\}$ is partitioned into $K$ clusters $\{\mathcal{C}_{\ell}\}_{\ell=1}^{K}$, and for each cluster fix a representative vector $v_{\ell}\in\mathbb{R}^{n}$ satisfying $\|v_{\ell}\|_{2}\leq\sqrt{n}$. Write $v_{r(j)}$ for the representative of the cluster containing $j$, and suppose the clusters are $\delta$-tight in the sense that
$\|X_{j}-v_{r(j)}\|_{2}\leq\delta\sqrt{n},\qquad\forall j,$